OpenStudy (anonymous):

FAN AND MEDAL PLEASE HELP f(x) = –16x2 + 24x + 16 Part A: What are the x-intercepts of the graph of f(x)? Show your work.

OpenStudy (anonymous):

@ganeshie8 @pooja195 @Kainui @kropot72 @Luigi0210 @Hero @abb0t @campbell_st

OpenStudy (anonymous):

@Hero please help

OpenStudy (anonymous):

ANYONE PLEASE HELP

OpenStudy (anonymous):

@triciaal PLEASE HELP

OpenStudy (anonymous):

PLZ HELP WOLF

OpenStudy (wolf1728):

When they ask for the x -intercepts, they are asking for the solution of the equation. Do you know how to solve it?

OpenStudy (anonymous):

0=-16x^2+24x+16 0 =-16x^2+32x-8x+16 0 =4x(-4x+8)+2(-4x+8) 0 = (4x+2)(-4x+8) 4x+2=0 and -4x+8=0 x=-1/2 and x=2 ??

OpenStudy (wolf1728):

Yes, you are correct!

OpenStudy (anonymous):

thank you!

OpenStudy (wolf1728):

u r welcome :-)

OpenStudy (triciaal):